OpenStudy (anonymous):

how do i do this? 8√5+3√5-√5

OpenStudy (amistre64):

let x = sqrt(5) 8x +3x -1x

OpenStudy (anonymous):

The square root of 5 would be like your variable and you would combine like terms.

OpenStudy (anonymous):

Since all of the roots are the same you just add and subtract the coefficient. 8+3-1=10. Therefore the answer is: \[10\sqrt{5}\]

OpenStudy (anonymous):

10√5

OpenStudy (anonymous):

so the root doesnt change?

OpenStudy (anonymous):

No it doesn't

OpenStudy (anonymous):

no because they're all the same root. It's like adding and subtracting fractions, if it has a common denominator you don't have to do much work.

OpenStudy (anonymous):

that makes alot of sense i hve a placement test for college this is on the practice test i dint knw hw to di now i do thank youuuu guys so much

OpenStudy (anonymous):

If you don't know radicals they'll place you in remedial. I had to teach this to my boyfriend he's stuck in MAT0028 -.-

OpenStudy (anonymous):

Make sure you know factoring, fractions and basic math too.

OpenStudy (anonymous):

loll aaahh nooo lol u cd retake it i think i have it in an hour :/ thank you :)

OpenStudy (anonymous):

I think you can retake it as long as its before school starts but you need to take your final scores to an adviser. Good luck :)

OpenStudy (anonymous):

thanks alot :) i really need it lol